OpenStudy (anonymous):

Find the value of x to the nearest tenth. 60.0 15.5 39.0 31.0

OpenStudy (blurbendy):

is there some sort of equation?

OpenStudy (whpalmer4):

use the property of similar triangles that corresponding sides are proportional. \[2x:62 = 39:(39+39) = 49:(49+49)\]Do the work carefully — there's a "trap" answer waiting for you that will be very tempting!

OpenStudy (anonymous):

thank @whpalmer4

OpenStudy (whpalmer4):

Did you get an answer yet?

OpenStudy (anonymous):

it is 39.0

OpenStudy (whpalmer4):

No :-( Maybe the ratio notation threw you off, I'll do it as fractions instead: \[\frac{2x}{62} = \frac{39}{39+39} = \frac{49}{49+49} = \frac{1}{2}\]

OpenStudy (anonymous):

so it is 15.5 then

OpenStudy (whpalmer4):

It couldn't be 39.0 — all of the other sides double when going from the small triangle to the big triangle, and 62 is not a doubling of 39.

OpenStudy (whpalmer4):

Yes, 15.5 is correct. The trap answer of 31 is also in the answer choices, to catch those who see that there's a ratio of 1:2 between the sides, but fail to notice that the problem calls the small side "2x" instead of "x" :-)

OpenStudy (anonymous):

sweet thanks you so much i was lost on how to set the problem up

OpenStudy (whpalmer4):

you're welcome.

OpenStudy (anonymous):

Which lines can you conclude are parallel given that M<7 + M<11 =180? Justify your conclusion with a theorem. Line a is parallel to line b by the Converse of the Alternate Interior Angles Theorem. Line c is parallel to line d by the Converse of the Same-Side Interior Angles Theorem. Line a is parallel to line b by the Converse of the Same-Side Interior Angles Theorem. Line c is parallel to line d by the Converse of the Alternate Interior Angles Theorem.